OpenStudy (anonymous):

Factor completely: 3x2 − x − 4 (3x − 1)(x + 4) (3x + 4)(x − 1) (3x − 2)(x + 2) (3x − 4)(x + 1)

OpenStudy (anonymous):

PLEASE!

OpenStudy (anonymous):

did you find the GCF (greatest common factor)?

OpenStudy (anonymous):

no i dont understand how to do that

OpenStudy (anonymous):

could you walk me through it

OpenStudy (anonymous):

okay well its a factoring problem this what you do, hold on

OpenStudy (anonymous):

let y= x^2 then think of it as: y^2-3y-4 then you have: (y+1)(y-4) so: (x^2 +1 ) (x^2 -4) now keep factoring: (x^2 +1)(x-2)(x+2) For high school math you'd stop here but, you can factor it more if you are using imagery numbers... (x-i)(x+i)(x-2)(x+2) where i = sqrt(-1) Hence the expression would have 4 roots when set equal to zero as expected.

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

yes thank you

OpenStudy (lncognlto):

The factors multiplied will be equal to the coefficient of x^2 and the constant multiplied, in this case 3 and -4, so -12. When added, they will be equal to the coefficient of the x term so -1 in this case. Thus the terms will be (3x − 4)(x + 1).

OpenStudy (lncognlto):

Sorry that was a little unclear. 3x2 − x − 4 = 3x2 + 3x - 4x -4 =3x(x+1) -4(x+1) =(3x-4)(x+1) The factors I was talking about were the coefficients of the x terms.